# TableBench: A Comprehensive and Complex Benchmark for Table Question Answering

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have markedly enhanced the interpretation and processing of tabular data, introducing previously unimaginable capabilities. Despite these achievements, LLMs still encounter significant challenges when applied in industrial scenarios, particularly due to the increased complexity of reasoning required with real-world tabular data, underscoring a notable disparity between academic benchmarks and practical applications. To address this discrepancy, we conduct a detailed investigation into the application of tabular data in industrial scenarios and propose a comprehensive and complex benchmark TableBench, including 18 fields within four major categories of table question answering (TableQA) capabilities. Furthermore, we introduce TableLLM, trained on our meticulously constructed training set TableInstruct, achieving comparable performance with GPT-3.5. Massive experiments conducted on TableBench indicate that both open-source and proprietary LLMs still have significant room for improvement to meet real-world demands, where the most advanced model, GPT-4, achieves only a modest score compared to humans. [^1] The instruction data, benchmark, and leaderboard are available at https://tablebench.github.io/..
1 Introduction
@@
Recent studies have shown the potential of large language models (LLMs) on tabular tasks such as table question answering (TableQA) (Zhu et al., 2021; Zhao et al., 2023; Hegselmann et al., 2023; Li et al., 2023b; Zhang et al., 2024b; Lu et al., 2024) by adopting in-context learning and structure-aware prompts (Singha et al., 2023), suggesting that a well-organized representation of tables improves the interpretation of tabular.  Tai et al. (2023) notes that eliciting a step-by-step reasoning process from LLMs enhances their ability to comprehend and respond to tabular data queries. Furthermore,  Zha et al. (2023) investigates the use of external interfaces for improved understanding of tabular data.
Figure 1: Typical challenges in TableBench: 1) Multi-hop fact-checking involves multiple steps to establish the relationship between facts across different years. 2) Multi-hop numerical reasoning requires calculating total operating expenses not directly provided in the table for specific years. 3) Trend forecasting involves estimating future data trends based on historical data analysis. 4) Chart generation necessitates executing program commands to create charts.
@@
Traditionally, adapting language models for tabular data processing entailed modifying their architectures with specialized features such as position embeddings and attention mechanisms to grasp tables’ structural nuances. However, the introduction of LLMs like GPT-4, GPT-3.5 (Brown et al., 2020; OpenAI, 2023), and PaLM2 (Anil et al., 2023) has heralded a new approach focused on the art of crafting precise, information-rich prompts that seamlessly integrate table data, coupled with leveraging external programming languages like SQL, Python, or other languages (Wang et al., 2023; Chai et al., 2024a), which facilitates more sophisticated chain-of-thought (Wei et al., 2022; Chai et al., 2024b) (CoT) reasoning processes across both proprietary and open-source LLM platforms, including Llama. Such advancements have propelled the fine-tuning of models for tabular data-specific tasks, showcased by initiatives like StructLM (Zhuang et al., 2024), enhancing capabilities in table structure recognition, fact verification, column type annotation, and beyond. However, the existing benchmark might not entirely resonate with the practical challenges, especially complex reasoning requirements encountered by professionals routinely navigating tabular data in real-world settings. Therefore, there is a huge need for creating a benchmark to bridge the gap between the industrial scenarios and the academic benchmark.
@@
To better evaluate the capability of LLMs in Table QA, we introduce TableBench, a comprehensive and complex benchmark covering 18 subcategories within four major categories of TableQA abilities. First, We systematically analyze real-world challenges related to table applications and define task complexity based on the required number of reasoning steps. Based on the analysis, we introduce a rigorous annotation workflow, integrating manual and automated methods, to construct TableBench. Subsequently, We create a massively TableQA instruction corpora TableInstruct, covering three distinct reasoning methods. Textual chain-of-thought (TCoT) utilizes a textual reasoning approach, employing a series of inferential steps to deduce the final answer. Symbolic chain-of-thought (SCoT) adopts symbolic reasoning steps, leveraging programming commands to iteratively simulate and refine results through a Think then Code process. Conversely, program-of-thought (PoT) generates executable code, using lines of code as reasoning steps within a programming environment to derive the final result. Based on open-source models and TableInstruct, we propose TableLLM as a strong baseline to explore the reasoning abilities of LLMs among tabular data, yielding comparable performance with GPT-3.5. Furthermore, we evaluate the performance of over 30 LLMs across these reasoning methods on TableBench, highlighting that both open-source and proprietary LLMs require substantial improvements to meet real-world demands. Notably, even the most advanced model, GPT-4, achieves only a modest score when compared to human performance.
@@
  • We systematically evaluate the interpretation and processing capabilities of 30+ models on our crafted TableBench and create a leaderboard to evaluate them on four main tasks. Notably, extensive experiments suggest that comprehensive and complex TableQA evaluation can realistically measure the gap between leading language models and human capabilities in real-world scenarios.
Figure 2: A comprehensive overview of the annotation framework.
2 Construction of TableBench
@@
2.1 Tabular Data Collection
We collect raw tabular data from existing datasets, including typical datasets such as WTQ (Pasupat and Liang, 2015), SQA (Iyyer et al., 2017), TabFact (Nan et al., 2022), FeTaQA (Nan et al., 2022), FinQA (Chen et al., 2021b), AIT-QA (Katsis et al., 2022), etc. We select tables based on topics and size, ensuring each contains at least 8 rows and 5 columns. We focus on tables with significant numerical values to emphasize numerical reasoning, thereby ensuring the complexity of the tabular data. Ultimately, we collect 3681 tables covering 20 major topics: finance, competition, sports, science, etc.
2.2 Question Annotation
2.3 Answer Annotation
2.4 Human Annotation
We paid all the annotators the equivalent of \$6 per question and provided them with a comfortable working environment, free meals, and souvenirs. We also provided the computer equipment and GPT-4 interface required for labeling. We labeled about 2,000 questions in total and employed them to check the quality of the questions/answers, and the total cost was about \$12,000 in US dollars. The annotators checked the derived tasks, including multilingual code explanation and code completion.
2.5 Quality Control
Iterative Annotation
We requested feedback from annotators on the task instructions and the website’s user experience, making iterative modifications to the annotation guidelines and website design based on this feedback.
Double-Blind Annotation
We employed a double-blind mechanism by assigning two independent annotators to each generated question or answer during the annotation process. These annotators worked independently, without any prior discussion. Following this, we conducted checks for annotation consistency. In cases of discrepancies, experts were invited to review the differing annotations, serving as guidelines to further refine our annotation standards, ensuring a comprehensive understanding of the constraints for specific categories of questions and the validity of computational results.
Quality Assessment Average Score Human Agreement(%) Randolph’s Kappa(%) / 95% CI
Table Complexity 3.94 79% 78% / [76%, 80%]
Question Complexity 4.12 88% 86% / [84%, 88%]
Objectivity of Answers 4.53 93% 92% / [89%, 94%]
Reasoning Steps 6.43 81% 78% / [74%, 82%]
Table 1: Quality assessment over 100 samples showcase the high consistency and complexity of TableBench
Quality Assessment
We invited five internal annotators to assess the quality of TableBench across four dimensions. Given the subjective nature of evaluating complexity and objectivity, annotators were asked to score based on their personal judgment. We designed a scoring system ranging from 1 to 5, providing reference examples and criteria for different scores. Reasoning steps follow the definition in 2.6. Additionally, we also reported an agreement percentage and Randolph’s Kappa (Randolph, 2005) (with 95% CI) in Table 1. The quantitative and qualitative analyses of the benchmark not only demonstrated its complexity but also highlighted the consistency of data annotations.
2.6 Dataset Statistic
@@
As illustrated in Figure 3, TableBench primarily comprises numerical tables, with the largest proportions being financial reports and data from competitive events.
Figure 3: Table Topic distribution
@@
Drawing from real-world scenarios and user demands for tabular data, we devise four primary question categories: fact-checking, numerical reasoning, data analysis, and visualization, encompassing 18 subcategories, thoroughly illustrating the various challenges encountered in TableQA scenarios. Compared with existing datasets, TableBench encompasses a broader spectrum of question categories as presented in Table 3, with a particular emphasis on data analysis and chart generation capabilities that are notably scarce in prior datasets.
Dataset Test Set Examples Fact Checking Numerical Reasoning Data Analysis Visulization
WTQ 4,344
SQA 3,012
TabFact 12,779
FeTaQA 2,000
FinQA 1,147
AIT-QA 515
WikiSQL 15,878
Spider 1,034
Bird 1,534
Text2Analysis 2,249
TableBench 886
Table 3: Comparison with existing datasets in categories.
Figure 4: Comparison with existing datasets in comlexity
@@
We define the complexity of the dataset by calculating the number of reasoning steps required to solve the problem. Figure 4 demonstrates that the overall complexity of TableBench is significantly higher than that of existing datasets, particularly in questions about data analysis and visualization.
3 TableLLM
3.1 Problem Definition
@@
Fact Checking Num-Reasoning Data Analysis Visualization Overall
DP TCoT SCoT PoT@$1$ DP TCoT SCoT PoT@$1$ DP TCoT SCoT PoT@$1$ DP@$1$ TCoT@$1$ SCoT@$1$ PoT@$1$ DP TCoT SCoT PoT@$1$
Human Performance 94.3 87.1 82.1 86.3 85.91
Open-source In Context Learning Methods
Llama2-7B-Chat 41.14 34.99 27.47 3.61 7.43 6.7 4.63 3.95 23.78 14.31 12.49 1.56 0.0 0.0 0.0 0.0 16.98 12.36 9.95 2.76
CodeLlama-7B-Instruct 42.5 33.06 12.34 19.44 6.1 5.43 2.99 13.31 24.88 16.16 17.06 1.79 0.0 0.0 0.0 0.0 17.01 12.3 9.28 8.85
Gemma-7B-Instruct 29.69 27.63 10.07 21.62 6.44 6.78 2.91 10.45 22.21 20.33 11.76 6.74 2.0 0.0 0.0 2.0 14.82 13.96 6.97 9.81
Mistral-7B-Instruct 50.15 50.45 40.56 6.25 9.32 8.73 5.77 2.6 24.55 21.99 21.12 1.19 0.0 0.0 0.0 0.0 19.15 17.86 15.11 2.35
Deepseek-Coder-7B-Instruct 21.6 22.92 27.48 48.98 4.75 6.45 5.61 34.66 22.17 18.73 20.72 18.17 14.0 8.0 18.0 18.0 13.82 13.1 14.58 28.89
CodeQwen1.5-7B-Chat 32.01 30.56 32.94 0.0 7.08 6.24 5.68 0.0 25.86 27.04 22.47 0.0 0.0 2.0 0.0 0.0 16.76 16.8 14.85 0.0
Qwen1.5-7B-Chat 26.47 56.08 53.53 39.23 8.04 11.3 10.99 20.4 24.13 24.77 22.96 7.66 0.0 0.0 0.0 0.0 15.84 20.7 19.65 16.29
Qwen2-7B-Instruct 53.76 57.7 57.52 0.0 15.75 16.09 16.65 0.76 21.51 24.02 21.5 0.38 0.0 0.0 4.0 2.0 21.23 22.77 22.26 0.6
StructLM-7B 34.64 47.72 64.06 13.54 5.79 9.55 19.97 11.48 14.37 19.59 23.83 4.38 2.0 0.0 0.0 0.0 12.06 17.06 25.21 8.3
MAP-Neo-7B-Instruct 22.43 32.7 33.22 0.0 7.17 7.23 6.46 0.0 18.21 21.85 14.38 0.44 0.0 0.0 0.0 4.0 12.66 15.26 12.03 0.4
Llama3-8B-Chat 70.79 38.32 72.53 13.94 18.55 22.02 17.33 19.5 28.93 30.15 30.75 9.31 0.0 0.0 0.0 10.0 27.28 25.71 27.59 14.43
Llama3.1-8B-Instruct 64.13 47.89 36.29 30.38 13.28 11.26 13.77 17.24 27.15 15.78 14.82 8.86 2.0 8.0 0.0 8.0 23.47 16.76 15.81 14.88
Llama2-13B-Chat 46.49 48.47 32.69 3.03 9.72 15.83 6.79 4.48 23.25 22.04 17.16 3.19 2.0 0.0 0.0 0.0 18.58 20.86 13.25 3.61
StructLM-13B 27.76 26.28 64.49 1.04 6.74 12.3 17.38 0.0 14.3 20.7 18.41 0.28 0.0 0.0 0.0 0.0 11.52 16.35 21.94 0.21
WizardLM-13B 52.35 53.93 46.01 8.33 11.48 13.79 16.52 14.79 25.5 22.61 20.16 3.73 0.0 0.0 0.0 4.0 20.8 20.75 20.23 9.12
Qwen1.5-14B-Chat 34.8 40.83 61.92 44.38 11.08 10.29 15.01 28.2 22.9 22.99 29.24 10.33 2.0 2.0 8.0 2.0 17.76 18.03 25.14 21.48
Qwen1.5-32B-Chat 42.32 64.99 67.86 49.01 15.6 19.13 21.15 34.01 21.46 24.27 28.29 17.43 4.0 4.0 8.0 8.0 20.21 25.38 28.3 27.79
Deepseek-Coder-33B-Instruct 28.38 48.27 54.34 33.12 5.09 9.41 12.69 32.6 10.13 9.09 21.7 19.97 8.0 0.0 0.0 24.0 9.74 13.01 19.92 27.2
CodeLlama-34B-Instruct 58.97 64.39 58.28 5.9 12.24 13.1 13.3 4.2 24.83 19.23 15.28 0.53 2.0 0.0 0.0 2.0 21.6 20.24 18.19 2.88
StructLM-34B 0.0 19.1 30.21 27.74 0.5 15.36 9.03 14.45 0.99 20.74 17.92 5.38 0.0 0.0 0.0 2.0 0.6 16.93 14.37 11.61
Mixtral-8x7B-Instruct 66.24 54.54 56.01 35.86 15.81 16.8 16.05 26.23 25.99 24.69 25.67 13.96 10.0 2.0 0.0 6.0 24.98 23.14 23.24 21.32
Qwen1.5-72B-Chat 75.34 71.27 67.03 33.16 21.1 19.01 16.68 20.85 28.02 26.63 27.33 13.03 0.0 2.0 8.0 14.0 28.45 26.66 25.8 18.65
Qwen2-72B-Instruct 71.11 72.5 71.13 56.37 27.32 36.97 31.81 41.33 30.9 32.2 31.85 22.36 10.0 20.0 14.0 12.0 32.52 38.13 35.14 33.91
Qwen1.5-110B-Chat 72.37 74.87 69.8 53.55 22.28 29.81 23.33 36.83 28.28 27.34 29.32 18.38 18.0 14.29 12.0 24.0 29.72 32.81 30.1 30.77
Llama3-70B-Chat 72.24 73.88 75.44 60.64 23.09 37.64 28.87 36.59 31.58 37.47 34.06 26.11 12.0 4.0 6.0 10.0 30.91 39.59 34.48 33.59
Llama3.1-70B-Instruct 72.31 76.32 77.65 59.05 27.6 44.89 38.93 34.04 31.14 35.88 33.87 23.15 22.0 26.0 6.0 34.0 33.63 43.85 39.22 32.52
Close-source In Context Learning Methods
GPT-3.5-Turbo 56.5 59.95 75.68 60.92 20.52 23.45 23.16 42.09 31.84 34.4 32.54 30.25 0.0 10.0 4.0 38.0 27.75 30.85 31.41 39.34
Qwen-Max 67.01 70.48 68.21 50.42 22.28 32.83 25.06 32.8 28.55 27.87 30.98 19.41 22.92 18.0 8.0 30.0 29.63 34.26 31.04 29.39
Yi-Large 70.26 71.41 66.08 13.19 25.01 40.18 23.2 15.25 29.74 29.22 22.59 5.97 37.5 26.0 26.0 6.0 32.43 38.57 27.82 10.9
GLM-4 70.63 67.93 73.59 31.49 24.99 34.01 26.18 25.46 31.19 30.47 31.54 25.34 6.0 8.0 14.0 22.0 31.23 34.8 32.76 25.92
Deepseek-Chat-V2 68.82 72.41 69.89 57.48 40.87 50.07 38.96 45.96 35.71 38.07 34.44 30.37 20.0 40.0 24.0 46.0 40.65 47.22 39.63 41.07
Deepseek-Coder-V2 70.25 74.58 34.9 62.91 31.48 48.46 18.23 47.08 31.07 33.52 13.09 31.62 26.0 40.0 26.0 44.0 35.21 45.02 18.5 42.67
GPT-4-Turbo 72.97 75.92 77.62 70.08 40.17 53.01 44.31 49.31 32.34 41.03 36.52 34.63 34.0 62.0 32.0 48.0 40.38 51.32 44.26 45.69
GPT-4o 74.29 72.63 71.01 62.31 42.73 54.46 42.2 47.83 34.36 38.9 34.65 30.03 38.0 56.0 38.0 54.0 42.73 50.39 42.22 42.92
Open-Source Fine-Tuning Methods
TableLLM-CodeQwen-7B 65.02 62.9 66.94 4.86 13.48 24.86 14.9 12.04 29.82 31.49 30.52 16.08 26.0 36.0 26.0 36.0 26.08 32.16 27.13 14.21
TableLLM-Deepseek-Coder-7B 67.32 69.23 63.15 7.12 17.14 35.87 21.2 14.61 29.86 31.75 29.62 21.04 26.0 36.0 18.0 30.0 27.98 37.76 28.8 17.19
TableLLM-Llama3.1-8B 63.68 68.15 65.17 25.67 18.01 30.51 17.86 28.64 28.32 33.47 30.2 19.77 24.0 24.0 18.0 44.0 27.19 35.29 27.7 25.76
TableLLM-Llama3-8B 66.03 62.13 64.46 15.07 16.59 29.42 16.73 12.68 28.02 30.21 29.45 17.53 28.0 26.0 20.0 28.0 26.93 32.89 26.97 15.83
TableLLM-Qwen2-7B 65.25 71.05 62.34 10.59 16.7 37.25 19.95 10.34 28.94 32.6 31.95 18.76 24.0 24.0 20.0 26.0 27.14 38.32 29.09 14.54
Table 4: The main results of advanced LLMs on TableBench are presented alongside human performance. Direct Prompting (DP) is a method that provides solutions directly without intermediate reasoning steps. PoT@$1$ refers to conducting code execution only once to derive the final answer. In the task of chart generation, all methods are required to generate code that is executed only once, evaluated with pass@$1$ as well.
3.2 Reasoning Methods
@@
In-context learning (Dong et al., 2022) (ICL) refers to strategies that optimize input for LLMs (($\mathcal{M}$)) to generate practical outputs with a task-specific instruction ($\mathcal{I}$) and a few output examples ($\mathcal{E}$). We introduce distinct reasoning methods to fully assess the reasoning capabilities of LLMs.
@@
TCoT (Wei et al., 2022) refers to a reasoning process in which LLMs incrementally derive a series of intermediate steps or sub-goals through textual prompts before generating the final answer. These intermediate steps constitute a "thought chain" that ultimately leads the model to the correct outcome. Formally, the method is:
@@
PoT (Chen et al., 2022) offers a novel approach to numerical reasoning tasks by distinctly delineating computation from reasoning. PoT decomposes the problem into programming commands $\mathcal{P}$ and utilizes a language interpreter, like Python, to compile and execute the resultant code. In contrast to SCoT, PoT enhances reasoning capabilities by actually executing generated code ($\mathcal{P}$) within a programming environment to output results, thereby implementing reasoning through structured code steps. The method can be formulated as:
@@
3.3 Supervised Fine-Tuning
The TableLLM training objective $\mathcal{L}_{all}$ of the Table QA instruction fine-tuning on TableInstruct can be described as:
@@ 3.3 Supervised Fine-Tuning
4 Experiments
4.1 Implementation Details
@@
Figure 5: The fine-grained categorical evaluation of proprietary models on TableBench with various reasoning methods.
4.2 LLMs
@@
We evaluate 34 models with sizes ranging from 7B to 110B parameters, including general/code LLMs, open-source/proprietary models, and SFT (Ouyang et al., 2022) models. For open-source LLMs, we evaluate on Llama2s (Touvron et al., 2023), Llama3s (Dubey et al., 2024), Llama3.1s, CodeLlamas (Roziere et al., 2023), CodeQwen1.5-7B-Chat, Qwen1.5s (Bai et al., 2023), Qwen2s (Yang et al., 2024a), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023), Deepseek-Coders (Guo et al., 2024), StructLMs (Zhuang et al., 2024), MAP-Neo-7B-Instruct (Zhang et al., 2024a), WizardLM-13B-V1.2 (Xu et al., 2023). For proprietary LLMs, we perform evaluation on GPTs (Brown et al., 2020; OpenAI, 2023) (GPT-3.5-Turbo, GPT4-Turbo, GPT4-o), Qwen-Max (Yang et al., 2024b), GLM-4 (GLM et al., 2024), Yi-Large (AI et al., 2024) and Deepseek models (Bi et al., 2024) (Chat-V2, Coder-V2). Furthermore, we finetune TableLLM based on CodeQwen-7B, DeepSeekCoder-7B, Llama3-8B, Llama3.1-8B, and Qwen2-7B to further explore the Table QA capabilities of LLMs.
4.3 Automatic Evaluation Metrics

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have markedly enhanced the interpretation and processing of tabular data, introducing previously unimaginable capabilities. Despite these achievements, LLMs still encounter significant challenges when applied in industrial scenarios, particularly due to the increased complexity of reasoning required with real-world tabular data, underscoring a notable disparity between academic benchmarks and practical applications. To address this discrepancy, we conduct a detailed investigation into the application of tabular data in industrial scenarios and propose a comprehensive and complex benchmark TableBench, including 18 fields within four major categories of table question answering (TableQA) capabilities. Furthermore, we introduce TableLLM, trained on our meticulously constructed training set TableInstruct, achieving comparable performance with GPT-3.5. Massive experiments conducted on TableBench indicate that both open-source and proprietary LLMs still have significant room for improvement to meet real-world demands, where the most advanced model, GPT-4, achieves only a modest score compared to humans.
Codehttps://github.com/TableBench/TableBench
Introduction
@@
Recent studies have shown the potential of large language models (LLMs) on tabular tasks such as table question answering (TableQA) (Zhu et al. 2021; Zhao et al. 2023; Hegselmann et al. 2023; Li et al. 2023b; Zhang et al. 2024b; Lu et al. 2024) by adopting in-context learning and structure-aware prompts (Singha et al. 2023), suggesting that a well-organized representation of tables improves the interpretation of tabular.  Tai et al. (2023) notes that eliciting a step-by-step reasoning process from LLMs enhances their ability to comprehend and respond to tabular data queries. Furthermore,  Zha et al. (2023) investigates the use of external interfaces for improved understanding of tabular data.
Figure 1: Typical challenges from four major categories (fact checking, numerical reasoning, data analysis, visualization) in TableBench.
Figure 2: A comprehensive overview of the annotation framework.
@@
Traditionally, adapting language models for tabular data processing entailed modifying their architectures with specialized features such as position embeddings and attention mechanisms to grasp structural nuances of tables. However, the introduction of LLMs like GPT-4, GPT-3.5 (Brown et al. 2020; OpenAI et al. 2024), and PaLM2 (Anil et al. 2023) has heralded a new approach focused on the art of crafting precise, information-rich prompts that seamlessly integrate table data, coupled with leveraging external programming languages like SQL, Python, or other languages (Wang et al. 2024; Chai et al. 2024), which facilitates more sophisticated chain-of-thought (Wei et al. 2022) (CoT) reasoning processes across both proprietary and open-source LLM platforms, including Llama. Such advancements have propelled the fine-tuning of models for tabular data-specific tasks, showcased by initiatives like StructLM (Zhuang et al. 2024), enhancing capabilities in table structure recognition, fact verification, column type annotation, and beyond. However, the existing benchmark might not entirely resonate with the practical challenges, especially complex reasoning requirements encountered by professionals routinely navigating tabular data in real-world settings. Therefore, there is a huge need for creating a benchmark to bridge the gap between the industrial scenarios and the academic benchmark.
@@
To better evaluate the capability of LLMs in Table QA, we introduce TableBench, a comprehensive and complex benchmark covering 18 subcategories within four major categories of TableQA abilities, as illustrated in Figure 1. First, We systematically analyze real-world challenges related to table applications and define task complexity based on the required number of reasoning steps. Based on the analysis, we introduce a rigorous annotation workflow, integrating manual and automated methods, to construct TableBench. Subsequently, We create a massively TableQA instruction corpora TableInstruct, covering three distinct reasoning methods. Textual chain-of-thought (TCoT) utilizes a textual reasoning approach, employing a series of inferential steps to deduce the final answer. Symbolic chain-of-thought (SCoT) adopts symbolic reasoning steps, leveraging programming commands to iteratively simulate and refine results through a Think then Code process. Conversely, program-of-thought (PoT) generates executable code, using lines of code as reasoning steps within a programming environment to derive the final result. Based on open-source models and TableInstruct, we propose TableLLM as a strong baseline to explore the reasoning abilities of LLMs among tabular data, yielding comparable performance with GPT-3.5. Furthermore, we evaluate the performance of over 30 LLMs across these reasoning methods on TableBench, highlighting that both open-source and proprietary LLMs require substantial improvements to meet real-world demands. Notably, even the most advanced model, GPT-4, achieves only a modest score when compared to human performance.
@@
  • We systematically evaluate the interpretation and processing capabilities of more than 30 models on our crafted TableBench and create a leaderboard to evaluate them on four main tasks. Notably, extensive experiments suggest that comprehensive and complex TableQA evaluation can realistically measure the gap between leading language models and human capabilities in real-world scenarios.
Construction of TableBench
@@
Tabular Data Collection
We collect raw tabular data from existing datasets, including typical datasets such as WTQ (Pasupat and Liang 2015), SQA (Iyyer, Yih, and Chang 2017), TabFact (Nan et al. 2022), FeTaQA (Nan et al. 2022), FinQA (Chen et al. 2021c), AIT-QA (Katsis et al. 2022), etc. To align closely with the ”reasoning complexity of questions” dimension in real-world tabular problems, we do not specifically design for the complexity of the tables themselves, such as structural complexity or large-sized tables. Instead, we adopt a moderate complexity in tabular data. We select tables based on topics and size, ensuring each contains at least 8 rows and 5 columns. We focus on tables with significant numerical values to emphasize numerical reasoning, thereby ensuring depth in numerical computation reasoning. Ultimately, we collect 3681 tables covering 20 major topics: finance, competition, sports, science, etc.
Figure 3: Topic distribution across tables
Dataset Fact Checking Numerical Reasoning Data Analysis Visulization
WTQ
SQA
TabFact
FeTaQA
FinQA
AIT-QA
WikiSQL
Spider
Bird
Text2Analysis
TableBench
Table 2: Comparison with existing datasets in categories.
(a) Reasoning steps of various question categories in TableBench
(b) Reasoning steps comparison across different datasets
Figure 4: Comparison with existing datasets in reasoning complexity.
Question Annotation
Answer Annotation
Dataset Statistic
@@
TableBench primarily consists of numerical tables, with the largest portions derived from financial reports and data from competitive events, as illustrated in Figure 3,
@@
Drawing from real-world scenarios and user demands for tabular data, we devise four primary question categories: fact-checking, numerical reasoning, data analysis, and visualization, encompassing 18 subcategories, thoroughly illustrating the various challenges encountered in TableQA scenarios, as shown in Table 1. Compared to existing datasets, TableBench covers a wider range of question types, as shown in Table 2, with a particular focus on data analysis and chart generation capabilities, which are notably lacking in previous datasets.
@@
We define the complexity of the dataset by calculating the number of reasoning steps required to solve the problem. Figure 4 illustrates that the overall complexity of the benchmark is significantly greater than that of existing datasets, especially concerning questions related to data analysis and visualization.
TableLLM
Problem Definition
@@
Reasoning Methods
@@
In-context learning (ICL) (Dong et al. 2022) refers to strategies that optimize input for LLMs ($\mathcal{M}$) to generate practical outputs with a task-specific instruction ($\mathcal{I}$) and a few output examples ($\mathcal{E}$). We introduce distinct reasoning methods to fully assess the reasoning capabilities of LLMs
@@
TCoT (Wei et al. 2022) refers to a reasoning process in which LLMs incrementally derive a series of intermediate steps or sub-goals through textual prompts before generating the final answer. These intermediate steps constitute a ”thought chain” that ultimately leads the model to the correct outcome. Formally, the method is:
@@
PoT (Chen et al. 2022) offers a novel approach to numerical reasoning tasks by distinctly delineating computation from reasoning. PoT decomposes the problem into programming commands $\mathcal{P}$ and utilizes a language interpreter, like Python, to compile and execute the resultant code. In contrast to SCoT, PoT enhances reasoning capabilities by actually executing generated code ($\mathcal{P}$) within a programming environment to output results, thereby implementing reasoning through structured code steps. The method can be formulated as:
@@
Supervised Fine-Tuning
We train TableLLM by fintuning all parameters of baseline LLMs to learn from the TableInstruct. The training objective $\mathcal{L}_{all}$ can be described as:
@@
Experiments
Implementation Details
@@
LLMs
@@
We evaluate 34 models with sizes ranging from 7B to 110B parameters, including general/code LLMs, open-source/proprietary models, and SFT (Ouyang et al. 2022) models. For open-source LLMs, we evaluate on Llama2s (Touvron et al. 2023), Llama3s (Grattafiori et al. 2024), Llama3.1s, CodeLlamas (Rozière et al. 2024), CodeQwen1.5-7B-Chat, Qwen1.5s (Bai et al. 2023), Qwen2s (Yang et al. 2024), Mistral-7B-Instruct-v0.2 (Jiang et al. 2023), Deepseek-Coders (Guo et al. 2024), StructLMs (Zhuang et al. 2024), MAP-Neo-7B-Instruct (Zhang et al. 2024a), WizardLM-13B-V1.2 (Xu et al. 2023). For proprietary LLMs, we perform evaluation on GPTs (Brown et al. 2020; OpenAI et al. 2024) (GPT-3.5-Turbo, GPT4-Turbo, GPT4-o), Qwen-Max (Yang et al. 2024), GLM-4 (GLM et al. 2024), Yi-Large (AI et al. 2024) and Deepseek models (DeepSeek-AI et al. 2024) (Chat-V2, Coder-V2). Furthermore, we finetune TableLLM based on CodeQwen-7B, DeepSeekCoder-7B, Llama3-8B, Llama3.1-8B, and Qwen2-7B to further explore the Table QA capabilities of LLMs.
Automatic Evaluation Metrics